# Evolution of LISA Observables for Binary Black Holes Lensed by an SMBH

### v1

@@ Abstract
Binary black holes (BBH) are expected to form and merge in active galactic nuclei (AGN), deep in the potential well of a supermassive black hole (SMBH), from populations that exist in a nuclear star cluster (NSC). Here we investigate the gravitational wave (GW) signature of a BBH lensed by a nearby SMBH. For a fiducial GW150914-like BBH orbiting close to a $10^{8}M_{\odot}$ SMBH located at $z=0.1$, the lensed GW signal varies in a predictable manner in and out of the LISA detectability band and across frequencies. The occurrence of such signatures has the potential to confound LISA global fit models if they are not modelled. Detection of these sources provide an independent measure of AGN inclination angles, along with detecting warping of the inner disk, and measuring the SMBH spin.
@@ 1 Introduction
The Laser Interferometer Space Antenna (LISA) is a European Space Agency (ESA) flagship mission to observe gravitational waves (GW) in the $f_{\rm GW}\sim 0.1-100$mHz frequency range (Amaro-Seoane et al., 2017, 2023), with an expected launch date in the mid-2030s. The mission will consist of 3 satellites forming a single triangular antenna. As a result, LISA will function as something of an ‘omni-scope’ with all simultaneously detected GW convolved together, thus requiring sophisticated data analysis techniques to extract and distinguish signals from various astrophysical sources to create a catalog which can be used by electromagnetic (EM) observers (e.g. Littenberg et al., 2013; Korol et al., 2017; Sberna et al., 2021). The challenging task of de-convolving these overlapping signals can be eased by making some assumptions about the nature of the observed astrophysical sources and their expected associated waveforms (Amaro-Seoane et al., 2023).
@@
Many GW sources are expected to emit in the LISA frequency range, including massive black hole binaries (MBHB–spanning total mass $\sim 10^{4-7}M_{\odot}$ (Sesana et al., 2011; Garg et al., 2024), white dwarf binaries (WDB) (Breivik et al., 2018; Littenberg et al., 2020), EMRIs & IMRIs (Kocsis et al., 2011; Amaro-Seoane, 2018; Derdzinski et al., 2021; Peng & Chen, 2023), and of most importance to this work, stellar origin binary black holes (BBH) in the early stages of their inspiral (Sesana, 2016a). These BBH are the same sources that are now routinely detected by higher frequency ground-based GW detectors (currently by the LIGO-Virgo-KAGRA [LVK] Consortium, e.g. (Abbott et al., 2023)). Ground-based GW detectors detect the final inspiral of such BBH, and with a lower strain sensitivity at higher GW frequencies, can detect many systems that LISA cannot observe. However, a number of higher mass, nearby, BBH systems should appear in the LISA band years to decades prior to their detection by ground-based GW observers (e.g. Sesana, 2016a; Wagg et al., 2022c; Sberna et al., 2022; McKernan et al., 2024). Such potential multi-band sources open up a wide range of interesting astrophysical questions, which could be answered using coordinated observing campaigns.
@@
The detectability of these BBH systems at wider separations in LISA rests on a number of assumptions: (1) $a_{\rm b,0}$, the semi-major axis of the binary at formation (if formed dynamically at smaller semi-major axis than appropriate to the LISA frequency range), (2) $\dot{a}_{\rm b}$, the speed of the semi-major axis decay (if purely GW driven or sped up through gas or dynamical interactions, e.g. $\ddot{a}_{\rm b}$), (3) ($a_{\rm b},e_{\rm b},i_{\rm b}$), the orbital parameters of the system appropriate to the LISA frequency band, including eccentricity around center of mass and inclination of orbit w.r.t. the observer’s sight-line, (4) $dN_{\rm b}/dt$, the rate of occurrence BBH (which is related to both their detection rate by ground-based detectors but also the factors listed above).
@@
Many authors have considered the LISA observability of BBH which may form from dynamical channels, notably in globular clusters (e.g. Kremer et al., 2018; Samsing & D’Orazio, 2018) or in Active Galactic Nucleus (AGN) accretion disks (e.g. McKernan et al., 2014; Toubiana et al., 2021; Sberna et al., 2022), as well as in isolated binary systems (e.g. Wagg et al., 2022c; van Zeist et al., 2023). The AGN channel of course rests on the assumption of a pre-existing nuclear star cluster (NSC) (e.g. Neumayer et al., 2020). The original GW detection of a BBH, GW150914 (Abbott et al., 2016) is often used as an example case, both because we know such systems exist, and if the system was evolving in vacuum, it would have been detectable by LISA (had LISA been observing) $\sim$ a decade before its merger could be seen from the ground (Sesana, 2016a). GW1501914-like BBH orbiting close enough to a supermassive black hole (SMBH) in an AGN that we would expect to see strong gravitational lensing effects imprinted on the LISA-observed waveform of the binary are only infrequently considered (e.g. Tamanini et al., 2020; Chen, 2021; Kuntz & Leyde, 2023; Zwick et al., 2023). While less likely, it is also possible for such a system to form in a gas-poor NSC, where it could experience similar lensing effects.
@@
Here we consider the simple case of a BBH orbiting the SMBH on a small, circular, co-planar orbit at 2 different radii within an AGN disk. We further assume the BBH itself has an eccentricity of zero ($e_{\rm b}=0$) around its own center of mass. We ignore gas or tidal effects on the BBH evolution, simply assuming vacuum evolution of $a_{\rm b}$ to examine the impact of lensing on the LISA-observed $f_{\rm GW}$ and strain ($h$) from such a source. We invoke AGN channel models for the formation of this hypothetical binary, finding that there is a potential to form a binary very close to the SMBH if multiple single black holes are efficiently delivered to the inner disk (e.g. Secunda et al., 2021). (Bellovary et al., 2016; Peng & Chen, 2021) have also shown that migration traps may exist at small disk radii, depending on the model (e.g. Sirko & Goodman, 2003), although see also (Grishin et al., 2024). For disks close to edge-on to the observer, BBH at small radii can experience extremely strong gravitational lensing effects from the SMBH, along with strong red-shifting due to the deep gravitational potential. The observed GW waveforms should depart substantially from vacuum expectation and will show strongly perturbed observed strain and frequencies, which will vary with the phase of the BBH orbit around the SMBH (Kuntz & Leyde, 2023). The BBH orbital timescale around the SMBH will be much smaller than the LISA mission lifetime, so these signatures varying periodically in ($f_{\rm GW},h$) must be accounted for if such a source is to be identified in the LISA data stream.
Figure 1: Illustration showing the physical setup for the environment in which we consider a GW150914 like binary to be in orbit around a $10^{8}~{}M_{\odot}$ SMBH. The center of mass of the binary is tracked in polar coordinates. The semi-major axis of the binary’s orbit is measured in gravitational radii $R_{g}$. The azimuthal angle $\varphi$ of the binary in the disk is measured from the axis extending to the observer. The inclination $\theta$ of the disk is measured from being face on to the observer, with the top of the disk moving away from the observer as the inclination increases.
@@ 2 Methods
We consider a binary system of the same description as GW150914 with masses of $36~{}M_{\odot}$ and $29~{}M_{\odot}$ placed at a distance of $410$ Mpc from a theoretical observer (Abbott et al., 2016). This system was chosen due to having an orbital frequency of $f_{orb}=0.8\times 10^{-3}$ Hz roughly five years before LIGO detected its coalescence, and would have been detectable by LISA had it been observing (Sesana, 2016b). We consider this binary in a circular, co-planer orbit, without perturbation, around an SMBH. The mass of the SMBH was chosen to be $10^{8}~{}M_{\odot}$ to artificially limit the effects of tidal disruption on the binary, if a full numerical simulation was run (see below).
Figure 1 shows the physical setup of our model and how the system is oriented with respect the observer. We assume the binary to be in a circularized orbit, so the position of the binary’s center of mass is tracked in polar coordinates. The orbital radius $r$ around the SMBH is measured in units of gravitational radii $R_{g}=GM_{\rm SMBH}/c^{2}$. The azimuthal angle is given as $\varphi$, where an observer exists along the axis extending from $\varphi=0^{\circ}$. The inclination angle of the disk $\theta$ is defined such that a disk with an inclination of $\theta=0^{\circ}$ would be face on to the observer. We can also specify the semi-major axis of the (presumed quasi-circular) BBH as $a$ (not pictured), as it evolves over time. Throughout the work, we assume an initial semi-major axis of $a=1\times 10^{-3}$ AU, which corresponds to an GW emission frequency of $f_{\rm{obs}}=1.6\times 10^{-3}$ Hz.
@@ 2.1 Ray tracing
The gravitational potential surrounding a central SMBH is typically described using either the Schwarzschild or the Kerr metric. Assuming that gravitational waves will propagate like photons through this metric (e.g. along null geodesics), we make use of preexisting ray-tracing solutions. In particular, we use Sim5111https://github.com/mbursa/sim5/ (Bursa, 2017) to trace geodesics from the observer to the mid-plane of an thin inclined accretion disk in the Kerr metric where the spin of the SMBH is set to 0.
@@
Certain regimes of parameter space require proper treatment of gravitational wave propagation using wave optics (D’Orazio & Loeb, 2020; Dalang et al., 2022). There are 3 major sources of interference which we should consider: scattering, magnification, and the modulation of the wave’s polarization. In the context of scattering, this may be defined as a three-body system (Pijnenburg et al., 2024). Wave optics become important in this type of system when the wavelength of the emitted GWs is longer than the Schwarzschild radius of the SMBH, which is the case when $f_{\rm{GW}}\geq 0.08\rm{Hz}\left(10^{5}M_{\odot}/M_{\rm{BH}}\right)$  (D’Orazio & Loeb, 2020). In this case, at any point of the orbit the GWs may be reflected off the SMBH and scattered to the observer, causing interference with the original waves (Dolan, 2008). For $f_{\rm{GW}}=1.6\times 10^{-3}\rm{Hz}$, this corresponds to SMBH with $M_{\rm{BH}}\leq 10^{6}M_{\odot}$, and our experiment is safely outside this range.
@@
Wave optics may also start becoming important in the context of gravitational lensing of the BBH system. The first order correction to geometric optics tells us that the GW’s tensorial distortion tensor from a lensing event is (Dalang et al., 2022):
@@
$\begin{split}F_{\mu\nu\alpha\beta}&=\left[m_{\mu}m_{\nu}-i\frac{4\Omega R_{\rm{s}}}{b^{2}}e^{2i\beta}n_{\mu}n_{\nu}\right]l_{\alpha}l_{\beta}\\ &\quad+\left[l_{\mu}l_{\nu}-i\frac{4\Omega R_{\rm{s}}}{b^{2}}e^{-2i\beta}n_{\mu}n_{\nu}\right]m_{\alpha}m_{\beta}\end{split}$ (1)
@@
with $\Omega\equiv 2\pi/\lambda$, and note that this is a first-order approximation for a point-like lens. The first set of terms in the brackets represent the undistorted tensor where geometric optics hold. The second term in each set of brackets represents deviations from the geometric optics regime, which scales $\propto R_{\rm{s}}\lambda^{-1}b^{-2}$ with the SMBH’s Schwarzschild radius, the inverse of the GW’s wavelength, and the inverse square of the impact parameter in gravitational radii. Within our experiment, the configuration most likely to lay outside the geometric optics regime is the inclined accretion disk at 88 with the BBH orbiting the SMBH at 20 $R_{\rm{g}}$. In this orbit when the BBH is farthest from the observer, the impact parameter becomes $b=20\cos{88^{\circ}}\sim 0.7$ (the impact parameter does not depend on the curvature of the metric). The redshifting of the GW by the SMBH’s gravitational well should be considered, which is approximately 0.5. We consider this by approximating the wavelengths observed by LISA as $\lambda_{\rm{emit}}=\lambda_{\rm{obs}}(1+z)^{-1}=cf_{\rm{obs}}^{-1}(1+z)^{-1}\approx 10^{-11}\rm{m}$. The distortion of the GW’s tensor is on the scale of $2\pi R_{\rm{s}}\lambda_{\rm{emit}}^{-1}b^{-2}\approx(10^{11}\rm{m})(10^{-11}\rm{m}^{-1})(2)\approx 2\approx b^{-2}$, meaning this is in the wave optics regime. This is an evolving system so we should determine the fraction of time we are within the wave optics regime. This may be approximated as the arc length of the orbit around the SMBH where the GW’s tensor is strongly affected by lensing. This may be estimated as when the correction terms in Eqn. (1) approach a considerable fraction of unity. We take this to be 0.1, defined as BBH positions with impact parameter $b<\sqrt{10}$ (only those positions behind the SMBH will experience lensing). For an orbit with radius $R_{\rm{orbit}}$ in units of gravitational radii, this is only a region within $\approx 9^{\circ}=\arccos\left(\sqrt{(1-10/R_{\rm{orbit}}^{2})/\sin^{2}(\theta)}\right)$ behind the SMBH. This limit is never reached in the configurations with $\theta=65^{\circ}$ or $R_{\rm{orbit}}=100R_{\rm{g}}$, and accounts for only $\sim$5of the orbit in the configuration $\theta=88^{\circ},R_{\rm{orbit}}=20R_{\rm{g}}$. Therefore, for our purposes of order of magnitude estimations across the 4 years of observations, we may use geometric optics to approximate the redshifting and amplification due to gravitational lensing.

### v2

@@ Abstract
Binary black holes (BBHs) are expected to form and merge in active galactic nuclei (AGN), deep in the potential well of a supermassive black hole (SMBH), from populations that exist in a nuclear star cluster (NSC). Here we investigate the gravitational wave (GW) signature of a BBH lensed by a nearby SMBH. For a fiducial GW150914-like BBH orbiting close to a $10^{8}M_{\odot}$ SMBH located at $z=0.1$, the lensed GW signal varies in a predictable manner in and out of the LISA detectability band and across frequencies. The occurrence of such signatures has the potential to confound LISA global fit models if they are not modelled. Detection of these sources provide an independent measure of AGN inclination angles, along with detecting warping of the inner disk, and measuring the SMBH spin.
@@ 1 Introduction
The Laser Interferometer Space Antenna (LISA) is a European Space Agency (ESA) flagship mission to observe gravitational waves (GW) in the $f_{\rm GW}\sim 0.1-100$mHz frequency range (Amaro-Seoane et al., 2017, 2023), with an expected launch date in the mid-2030s. The mission will consist of three satellites forming a single triangular antenna. As a result, LISA will function as something of an ‘omni-scope’ with all simultaneously detected GW convolved together, thus requiring sophisticated data analysis techniques to extract and distinguish signals from various astrophysical sources to create a catalog which can be used by electromagnetic (EM) observers (e.g. Littenberg et al., 2013; Korol et al., 2017; Sberna et al., 2021). The challenging task of de-convolving these overlapping signals can be eased by making some assumptions about the nature of the observed astrophysical sources and their expected associated waveforms (Amaro-Seoane et al., 2023).
@@
Many GW sources are expected to emit in the LISA frequency range, including massive black hole binaries (MBHB–spanning total mass $\sim 10^{4-7}M_{\odot}$ (Sesana et al., 2011; Garg et al., 2024), white dwarf binaries (WDB) (Breivik et al., 2018; Littenberg et al., 2020), extreme mass ratio inspirals (EMRIs) & intermediate mass ratio inspirals (IMRIs) (Kocsis et al., 2011; Amaro-Seoane, 2018; Derdzinski et al., 2021; Peng & Chen, 2023), and of most importance to this work, stellar-mass binary black holes (BBHs) in the early stages of their inspiral (Sesana, 2016a). These BBHs are the same sources that are now routinely detected by higher frequency ground-based GW detectors (currently by the LIGO-Virgo-KAGRA [LVK] Consortium, e.g. (Abbott et al., 2023)). At the time of coalescence, BBHs are emitting at a strain sensitivity and GW frequency only detectable by ground-based GW detectors. However, a number of higher mass, nearby, BBH systems should appear in the LISA band years to decades prior to their detection by ground-based GW observers (e.g. Sesana, 2016a; Wagg et al., 2022c; Sberna et al., 2022; McKernan et al., 2024). Such potential multi-band sources open up a wide range of interesting astrophysical questions, which could be answered using coordinated observing campaigns.
@@
The following conventions are used to describe the parameters for a BBH system detectable by LISA: (1) $a_{\rm b,0}$, the semi-major axis of the binary at formation (if formed dynamically at smaller semi-major axis than appropriate to the LISA frequency range), (2) $\dot{a}_{\rm b}$, the decay rate of the semi-major axis (if purely GW driven or sped up through gas or dynamical interactions, e.g. $\ddot{a}_{\rm b}$), (3) ($a_{\rm b},e_{\rm b},i_{\rm b}$), the orbital parameters of the system appropriate to the LISA frequency band, including eccentricity around center of mass and inclination of orbit w.r.t. the observer’s sight-line, (4) $dN_{\rm b}/dt$, the rate of occurrence of BBHs (which is related to both their detection rate by ground-based detectors but also the factors listed above).
@@
Many authors have considered the LISA observability of BBHs which may form from dynamical channels, notably in globular clusters (e.g. Kremer et al., 2018; Samsing & D’Orazio, 2018) or in Active Galactic Nucleus (AGN) accretion disks (e.g. McKernan et al., 2014; Toubiana et al., 2021; Sberna et al., 2022). The AGN channel of course rests on the assumption of a pre-existing nuclear star cluster (NSC) (e.g. Neumayer et al., 2020). Isolated binary channels are also sources of LISA detectable BBHs (e.g. Wagg et al., 2022c; van Zeist et al., 2023), forming from stable mass transfer (e.g. Gallegos-Garcia_2021), chemically homogeneous stellar binaries (e.g. 10.1093/mnras/stw379), and the stellar common envelope (e.g. Dominik2012ApJ). The original GW detection of a BBH, GW150914 (Abbott et al., 2016) is often used as an example case, both because we know such systems exist, and if the system was evolving in vacuum, it would have been detectable by LISA (had LISA been observing) $\sim$ a decade before its merger could be seen from the ground (Sesana, 2016a). If a GW150914-like BBH is orbiting close to a supermassive black hole (SMBH) in an AGN, we expect strong gravitational lensing effects. These effects have been infrequently considered in previous literature (e.g. Tamanini et al., 2020; Chen, 2021; Kuntz & Leyde, 2023; Zwick et al., 2023). Furthermore, it would not be unusual for such a binary to form in an AGN (McKernan et al., 2024).In the LVK band, the rate of occurrence for BBH mergers near the last migration trap is analytically estimated to be $\sim 0.4~\text{Gpc}^{-3}\text{yr}^{-1}$ for a $3\times 10^{8}M_{\odot}$ SMBH, where the AGN lifespan is $10^{7}~\text{yrs}$ (Peng & Chen, 2021). We will analyze the uncertainties in this rate in our discussion section. While less likely, it is also possible for such a system to form in a gas-poor NSC, where it could experience similar lensing effects.
@@
Here we consider the simple case of a BBH orbiting the SMBH on a small, circular, co-planar orbit at 2 different radii within an AGN disk. We further assume the BBH itself has an eccentricity of zero ($e_{\rm b}=0$) with respect to its own center of mass, consistent with the observed eccentricity of GW150914. We ignore gas and tidal effects on the BBH’s evolution, simply assuming vacuum evolution of $a_{\rm b}$ to examine the impact of lensing on the LISA-observed $f_{\rm GW}$ and strain ($h$) from such a source. We invoke AGN channel models for the formation of this hypothetical binary, finding that there is a potential to form a binary very close to the SMBH if multiple single black holes are efficiently delivered to the inner disk (e.g. Secunda et al., 2021). Bellovary et al. (2016); Peng & Chen (2021) have also shown that migration traps may exist at small disk radii, depending on the model (e.g. Sirko & Goodman, 2003), although also see Grishin et al. (2024). For disks close to edge-on to the observer, BBHs at small radii can experience extremely strong gravitational lensing effects from the SMBH, along with strong red-shifting due to the deep gravitational potential. The observed GW waveforms should depart substantially from vacuum expectation and will show strongly perturbed observed strain and frequencies, which will vary with the phase of the BBH’s orbit around the SMBH (Kuntz & Leyde, 2023). The BBH’s orbital timescale around the SMBH will be much smaller than the LISA mission lifetime, so these signatures varying periodically in ($f_{\rm GW},h$) must be accounted for if such a source is to be identified in the LISA data stream.
Figure 1: Illustration showing the physical setup for the environment in which we consider a GW150914 like binary to be in orbit around a $10^{8}~M_{\odot}$ SMBH. The center of mass of the binary is tracked in polar coordinates. The semi-major axis of the binary’s orbit is measured in gravitational radii $R_{g}$. The azimuthal angle $\varphi$ of the binary in the disk is measured from the axis extending to the observer. The inclination $\theta$ of the disk is measured from being face-on to the observer, with the top of the disk moving away from the observer as the inclination increases.
@@ 2 Methods
We consider a binary system of the same description as GW150914 with masses of $36~M_{\odot}$ and $29~M_{\odot}$ placed at a distance of $410$ Mpc from a theoretical observer (Abbott et al., 2016). Along with the BBH being nearly circular, previous studies focusing on the possible multi-band detectability of GW150914 make it an obvious fiducial system to base our tests on. The system would have had an orbital frequency of $f_{orb}=0.8\times 10^{-3}$ Hz roughly 5 years before LIGO detected its coalescence, and detectable by LISA had it been observing (Sesana, 2016a). We consider this binary in a circular, co-planar orbit, without perturbation, around an SMBH. The mass of the SMBH was chosen to be $10^{8}~M_{\odot}$ to artificially limit the effects of tidal disruption on the binary, if a full numerical simulation was run (see below).
Figure 1 shows the physical setup of our model and how the system is oriented with respect the observer. We assume the binary to be in a circularized orbit, so the position of the binary’s center of mass is tracked in polar coordinates. The orbital radius $r$ around the SMBH is measured in units of gravitational radii $R_{g}=GM_{\rm SMBH}/c^{2}$. The azimuthal angle is given as $\varphi$, where an observer exists along the axis extending from $\varphi=0^{\circ}$. The inclination angle of the disk $\theta$ is defined such that a disk with an inclination of $\theta=0^{\circ}$ would be face-on to the observer. We pick inclination angles of $\theta=88^{\circ}$ and $\theta=65^{\circ}$, so we can explore one highly inclined case and one less inclined case, respectively. We further discuss the range of inclination angles, for which our results are valid, in section 4.1 below. We can also specify the semi-major axis of the (presumed quasi-circular) BBH as $a_{\textrm{b}}$ (not pictured), as it evolves over time. Throughout the work, we assume an initial semi-major axis of $a_{\textrm{b}}=1\times 10^{-3}$ AU, which corresponds to an GW emission frequency of $f_{\rm{obs}}=1.6\times 10^{-3}$ Hz.
@@ 2.1 Ray tracing
The gravitational potential surrounding a central SMBH is typically described using either the Schwarzschild or the Kerr metric. Assuming that gravitational waves propagate along null geodesics (e.g., like photons), we make use of preexisting ray-tracing solutions. In particular, we use Sim5111https://github.com/mbursa/sim5/ (Bursa, 2017) to trace geodesics from the observer to the mid-plane of an thin inclined accretion disk. Sim5 assumes a Kerr metric; however, for simplicity, we set the spin of the SMBH to 0. This results in a Schwarzschild metric, meaning our results only apply to cases where the SMBH is not spinning. Since a spinning SMBH is well motivated, it would be worth repeating our investigation for a variety of SMBH spins in future work.
@@
Certain regimes of parameter space require proper treatment of gravitational wave propagation using wave optics (D’Orazio & Loeb, 2020; Dalang et al., 2022). There are three major sources of interference which we should consider: scattering, magnification, and the modulation of the wave’s polarization. If we consider the scattering of gravitational waves by a third body (e.g. the SMBH), we must determine how strong the scattered waves appear to the observer and if they may considerably interfere with the original source (Dolan, 2008; Pijnenburg et al., 2024). As described in D’Orazio & Loeb (2020), wave optics become important in this type of system when the wavelength of the emitted GWs is longer than the Schwarzschild radius of the third black hole, which is the case when $f_{\rm{GW}}\geq 0.08\rm{Hz}\left(10^{5}M_{\odot}/M_{\rm{BH}}\right)$. This effect does not require traditional lensing, and will occur at any point of the orbit the BBH. For $f_{\rm{GW}}=1.6\times 10^{-3}\rm{Hz}$, this corresponds to SMBH with $M_{\rm{BH}}\leq 10^{6}M_{\odot}$, and our experiment is safely outside this range due to the large mass of the SMBH.
@@
Wave optics may also start becoming important in the context of gravitational lensing of the BBH system. To estimate the first order correction to geometric optics, we follow section 3 and 4 of Dalang et al. (2022) to determine the magnitude of distortions of the gravitational wave [^2] Note that this is a first-order approximation for a point-like lens.:
@@
$\begin{split}F_{\mu\nu\alpha\beta}&=\left[m_{\mu}m_{\nu}-i\frac{4\Omega R_{\rm{s}}}{b^{2}}e^{2i\delta}n_{\mu}n_{\nu}\right]l_{\alpha}l_{\beta}\\ &\quad+\left[l_{\mu}l_{\nu}-i\frac{4\Omega R_{\rm{s}}}{b^{2}}e^{-2i\delta}n_{\mu}n_{\nu}\right]m_{\alpha}m_{\beta},\end{split}$ (1)
@@
where $i\equiv\sqrt{-1}$ the imaginary unit, $\Omega\equiv 2\pi/\lambda_{\rm{emit}}$ the energy of the wave, $\lambda_{\rm{emit}}$ is the emitted wavelength from the GW source, $R_{\rm{s}}$ is the Schwarzschild radius of the SMBH, $b$ is the impact parameter often measured in units of $R_{\rm{g}}$, $\delta$ is the angle between the plus polarization and the impact position in the lens plane, and $n_{\mu},m_{\mu},\text{ and }l_{\mu}$ are basis vectors such that in the unperturbed metric: $\bar{n}^{\mu}\propto(1,0,0,-1),\bar{m}^{\mu}\propto(0,1,i,0),\text{and }\bar{l}^{\mu}=\bar{m}^{*\mu}$. To first order, the perturbed components become: $n^{\mu}\propto(1,0,0,-1),m^{\mu}\propto(-R_{\rm{s}}b^{-1}e^{i\delta},1,i,R_{\rm{s}}b^{-1}e^{i\delta}),l^{\mu}\propto(-R_{\rm{s}}b^{-1}e^{-i\delta},1,-i,R_{\rm{s}}b^{-1}e^{-i\delta})$. In equation (1), the first set of terms in the brackets represent the undistorted tensor where geometric optics hold. Importantly, the second term in each set of brackets represents deviations from the geometric optics regime, which scales $\propto R_{\rm{s}}\lambda_{emit}^{-1}b^{-2}$. Within our experiment, the configuration most likely to lay outside the geometric optics regime is the inclined accretion disk at 88 with the BBH orbiting the SMBH at 20 $R_{\rm{g}}$. In this orbit, when the BBH is farthest from the observer, the impact parameter becomes $b=20R_{\rm{g}}\cos{88^{\circ}}\sim 0.7R_{\rm{g}}$ (the impact parameter does not depend on the curvature of the metric). The redshifting of the GW by the SMBH’s gravitational well and orbital motion should be considered, which is approximately 0.5 when the BBH is 20 $R_{\rm{g}}$ from the SMBH and appears to be receding from the observer (e.g. see Figure 4). We consider this by approximating the wavelengths observed by LISA as $\lambda_{\rm{emit}}=\lambda_{\rm{obs}}(1+z)^{-1}=cf_{\rm{obs}}^{-1}(1+z)^{-1}\approx 10^{-11}\rm{m}$. From equation (1), we determine that the magnitude of distortion is approximately $2\pi R_{\rm{s}}\lambda_{\rm{emit}}^{-1}b^{-2}$, and with relevant values becomes $\approx(10^{11}\rm{m})(10^{-11}\rm{m}^{-1})(2)\approx 2$. We note that this reduces to distortions that are $\approx b^{-2}$, meaning this is in the wave optics regime and our geometric solution does not truly represent the expected modulation of signal. However, this is an evolving system so we should determine if there is a significant amount of time that we are within the wave optics regime. This may be approximated as the arc length of the orbit around the SMBH where the GW’s tensor is strongly affected by lensing. This may be estimated as when the correction terms in equation (1) approach a considerable fraction of unity. We define this threshold to be 0.1, defined as the BBH’s position with an impact parameter $b<\sqrt{10}R_{\rm{g}}$, representing a 10% deviation from geometric optics. For an orbit with radius $R_{\rm{orbit}}$ in units of gravitational radii, this is only a region within $\approx 9^{\circ}=\arccos\left(\sqrt{(1-10R^{2}_{\rm{g}}/R_{\rm{orbit}}^{2})/\sin^{2}(\theta)}\right)$ behind the SMBH. This limit is never reached in the configurations with $\theta=65^{\circ}$ or $R_{\rm{orbit}}=100R_{\rm{g}}$, and accounts for only $\sim$5% of the orbit in the configuration $\theta=88^{\circ},R_{\rm{orbit}}=20R_{\rm{g}}$. Therefore, for our purposes of order of magnitude estimations across the 4 years of observations, we may use geometric optics to approximate the redshifting and amplification due to gravitational lensing.